### v3

# LTM: Scalable and Black-box Similarity-based Test Suite Minimization based on Language Models

Abstract
Test suites tend to grow when software evolves, making it often infeasible to execute all test cases with the allocated testing budgets, especially for large software systems. Test suite minimization (TSM) is employed to improve the efficiency of software testing by removing redundant test cases, thus reducing testing time and resources, while maintaining the fault detection capability of the test suite. Most existing TSM approaches rely on code coverage (white-box) or model-based features, which are not always available to test engineers. Recent TSM approaches that rely only on test code (black-box) have been proposed, such as ATM and FAST-R. To address the scalability, we propose LTM (Language model-based Test suite Minimization), a novel, scalable, and black-box similarity-based TSM approach based on large language models (LLMs), which is the first application of LLMs in the context of TSM. To support similarity measurement for test code embeddings, we investigate five pre-trained language models: CodeBERT, GraphCodeBERT, UniXcoder, StarEncoder, and CodeLlama, on which we compute two similarity measures: Cosine Similarity and Euclidean Distance. Our goal is to find similarity measures that are not only computationally more efficient but can also better guide a Genetic Algorithm (GA) to search for optimal minimized test suites, thus reducing the overall search time. Experimental results show that the best configuration of LTM (UniXcoder/Cosine) outperforms ATM in three aspects: (a) achieving a slightly greater saving rate of testing time ($41.72\%$ versus $40.29\%$, on average); (b) attaining a significantly higher fault detection rate ($0.84$ versus $0.81$, on average); and, most importantly, (c) minimizing test suites nearly five times faster on average, with higher gains for larger test suites and systems, thus achieving much higher scalability.
Index Terms:
Test suite minimization, Test suite reduction, Pre-trained language models, Genetic algorithm, Black-box testing
1 Introduction
Software testing is an essential activity that ensures high-quality software systems by detecting faults in software releases [1]. When test suites are large, especially for large code bases, the scalability of running all test cases in a test suite becomes quickly a critical issue, as time and resources are limited for testing in industrial contexts [2]. The problem is particularly acute as test suites tend to increase in size over time with software evolution, making it impossible to run all test cases for every code change with allocated testing budgets [1]. To address this issue, test suite minimization (TSM) is widely employed to make the testing process more efficient while maintaining its fault detection capability [1]. In contrast to test case selection [3], which selects a subset of test cases that are relevant to code changes, and test case prioritization [3], which ranks test cases based on certain characteristics, such as fault detection capability, the objective of test suite minimization is to eliminate redundant or similar test cases that are unlikely to detect different faults [3, 1].
Though there exist various TSM approaches, most of them rely on either code coverage information (white-box) [3], which requires access to software production code, or model-based features, such as transition coverage for UML state machine-based testing [2]. This information is not always accessible by test engineers and can be challenging to collect [4, 5]. Further, collecting code coverage information is expensive [6, 7] and can lead to scalability issues for large software systems. Cruciani et al. [6] and Pan et al. [8] addressed this issue by proposing TSM approaches that rely only on the source code of test cases (black-box). FAST-R, which is based on clustering algorithms, has been shown to be much more efficient than white-box approaches, but achieved relatively low fault detection rates. ATM, which is based on test case similarity and evolutionary search, achieved higher fault detection rates within practically acceptable, yet longer time. Since test suite minimization is performed on major software releases [9, 8], as opposed to test case selection and prioritization which are performed for every code change, ATM provides a better trade-off compared to FAST-R in many practical situations. However, ATM still presents scalability issues for very large software systems, since its total minimization time (the sum of preparation time and search time) increases rapidly with test suite size, represented by the number of test cases per version. Our analysis of the total minimization time of ATM reveals that similarity measures play a major role in limiting its scalability, due to the fact that (1) computing tree-based similarity is expensive, taking up to $41.20\%$ of the total minimization time, and (2) similarity measures impact the search convergence and speed. This motivated us to investigate similarity measures that are both more efficient to calculate and more informative to guide the search.
Language models pre-trained on programming languages convert test code into vector-based embeddings. This enables vector-based similarity measurement, which facilitates implementation optimizations and results in much higher computational efficiency than tree-based similarity measurement that relies on traversing AST trees. Moreover, these language models are pre-trained on large source code corpora with various code understanding and generation tasks, thus generating embeddings that capture informative syntactic and contextual details from test code. This suggests that such embeddings with vector-based similarity might be potentially more informative than tree-based similarity in guiding evolutionary search. Therefore, in this paper, we propose LTM (Language model-based Test suite Minimization), a scalable, black-box similarity-based TSM approach that is based on pre-trained Large Language Models (LLMs) and vector-based similarity measures, making it the first application of LLMs in the context of TSM, to the best of our knowledge. LTM uses the source code of test cases (Java test methods), without requiring any preprocessing, as input to five alternative pre-trained language models, namely CodeBERT [10], GraphCodeBERT [11],UniXcoder [12], StarEncoder [13], and CodeLlama [14], in order to extract test method embeddings. Considering that test suite diversity was reported to have a positive correlation with fault detection [2, 15, 16], LTM employs two similarity measures: Cosine similarity and Euclidean distance, to calculate the similarity between the extracted embeddings. Using the calculated similarity values, LTM employs a Genetic Algorithm (GA) to minimize test suites, which searches for the most optimal subset of a test suite, for a given testing budget. We evaluated LTM with the same dataset used to evaluate ATM (Defects4J), followed a similar experimental design, and used the same evaluation metrics of effectiveness and efficiency: Fault Detection Rate (FDR) and Minimization Time (MT), respectively. In addition, considering the potential variation in test case execution times, assessing minimized test suites based on their number of test cases only might not always be accurate [1]. Therefore, we extended our evaluation of minimized test suites using an additional metric: Time Saving Rate (TSR). Moreover, we optimized GA by utilizing a more efficient data structure to accelerate fitness calculation and enhance memory usage, which led to a $273$-fold reduction in minimization time, without requiring any additional computation resources. Then, we identified the best configuration of LTM by considering both effectiveness and efficiency, among all its alternatives, and compared it to the two best ATM configurations. Finally, we expanded our experiments by running LTM on a much larger project that ATM could not handle, to further assess the scalability of our approach.
Specifically, we address the following research questions.
  • RQ1: How does LTM perform for test suite minimization under different configurations? LTM achieves high FDR results (an overall average FDR of $0.79$ across configurations) for a $50\%$ minimization budget (i.e., the percentage of test cases retained in the minimized test suite). The best configuration of LTM is UniXcoder using Cosine similarity when considering both effectiveness ($0.84$ FDR on average) and efficiency ($0.82$ min on average), which also achieves a greater time saving rate (an average TSR of $41.72\%$). For the large project, $Closure$, UniXcoder using Cosine Similarity takes only $17.90$ min in terms of MT and achieves an FDR of $0.79$, while saving $52.55\%$ of testing time.
  • RQ2: How does LTM compare to ATM? The best configuration of LTM outperforms ATM by achieving significantly better FDR ($0.84$ versus $0.81$, on average), and more importantly, running much faster than ATM ($0.82$ min versus $4.06$ min, on average), in terms of both preparation time (up to two orders of magnitude faster) and search time (up to one order of magnitude faster). The latter is particularly important on typically large industrial systems and test suites where such differences practically matter.
To summarize, the contributions of this work are as follows:
  • We propose a novel black-box TSM approach (LTM) that relies, for the first time, on LLMs and two distance functions based on the generated embeddings. We investigate and conduct a comprehensive comparison among five recent language models with various model architectures, parameter sizes, inputs, and tasks used for pre-training.
  • We optimize the search process of LTM by utilizing a more efficient data structure for fitness calculation, which in turn reduced the search time by $273$ folds.
  • We conduct a thorough comparative analysis of the efficiency and effectiveness, as well as the achieved saving in testing time, of black-box TSM approaches based on a large-scale test suite minimization experiments involving $17$ Java projects with $835$ versions, thus yielding valuable insights into the relative performance of alternatives. Overall, the experiments took around three months in calendar time corresponding to 6 years of computation on a cluster with $83,216$ available CPU cores.
  • We analyze the minimization time of LTM and show that it is much more scalable than the state-of-the-art (SOTA) approaches by running five times faster on average—with even higher gains for larger systems and test suites typically encountered in practice—while achieving significantly higher fault detection rates, as a result of using LLMs for embeddings.
The rest of this paper is organized as follows. Section 2 reviews related work and motivates our research. Section 3 describes our test suite minimization approach. Section 4 validates our approach, reports the experimental design and results, and discusses the implications in practice. Section 5 discusses the validity threats to our results. Section 6 draws conclusions and suggests future work.
2 Related Work
Test suite minimization (TSM) aims at improving the efficiency of software testing by removing redundant test cases, thus reducing testing time and resources while maintaining the effectiveness of the test suite (i.e., fault detection capability) [1]. There are various approaches that have been proposed to address it, including (a) greedy heuristics-based approaches [17, 9], which select test cases iteratively based on code coverage information, (b) clustering-based approaches [18, 19], which group test cases based on the similarity of their coverage information, and (c) search-based approaches [20, 2], which employ evolutionary search to find an optimal subset of the test suite using model-based features. Most of these approaches utilize code coverage information (white-box) [3] (i.e., requires access to the system production code) or model-based features, which are not always available to test engineers [8, 5]. Moreover, collecting code coverage information can result in up to $30\%$ time overhead [6, 7], making it not scalable for large test suites and systems.
Viggiato et al. [21] investigate the similarity of test cases that are written in natural language. Though their objective is to identify redundancy among test cases, as opposed to TSM within a budget, their similarity measures can priori be further used for this purpose. They convert test cases into vector-based representations using five text embedding techniques, i.e., Word2Vec [22], BERT [23], Sentence-BERT [24], Universal Sentence Encoder [25] and TF-IDF (Term Frequency–Inverse Document Frequency) [26]. They then identify similar test cases utilizing clustering algorithm to group similar test steps based on two different similarity metrics (Word Mover’s Distance (WMD) [27] and Cosine Similarity). The results show that their approach achieves an F-Score of $83.47\%$ for identifying similar test cases. In summary, different from our work, their focus is on identifying redundant test cases written in natural language, as opposed to code, and they do not aim to optimize a test suite within a budget.
Philip et al. [28] proposed a black-box approach, called FastLane, which relies on test and commit logs containing historical information, to predict test case outcomes (i.e., pass or fail) using logistic regression to skip their execution. Using such information, FastLane reduces testing time and resources by running only a subset of test cases instead of the whole test suite. Results showed that FastLane reached $99.99\%$ in terms of test outcome accuracy and saved up to $18.04\%$ of testing time. However, it requires historical information about multiple runs of test cases, thus making FastLane inapplicable for new or recent test cases. Moreover, FastLane cannot be easily adapted to given minimization budgets (i.e., a predefined percentage of test cases to be executed). Arrieta et al. [5] relied on test case inputs and outputs for simulation-based testing and employed Non-Dominated Sorting Genetic Algorithm II (NSGA-II) [29] to find an optimal subset of test cases as a minimized test suite. Chang et al. [30] detected redundancy in test cases written in natural language using word embedding techniques and Cosine similarity. However, such information is not accessible in many contexts, thus rendering such black-box approaches not applicable.
Cruciani et al. [6] proposed a black-box approach, called FAST-R, that relies solely on the source code of test cases. FAST-R converts test code into vectors using a term frequency model [31] and then employs a random projection technique [32] to reduce the dimensionality of vectors. Based on these vectors, clustering-based algorithms were performed with the centroids of clusters selected as a minimized test suite. Results showed that, compared to white-box approaches, FAST-R was much more efficient in terms of total minimization time while achieving comparable effectiveness in terms of fault detection capability. However, FAST-R achieved relatively lower fault detection capability for Java projects, with median fault detection rates ranging from $0.18$ to $0.22$ for minimization budgets ranging from $1\%$ to $30\%$. With such low effectiveness, FAST-R is therefore not a viable option in many practical contexts.
To achieve a better trade-off between the effectiveness and efficiency of test suite minimization, Pan et al. [8] proposed a black-box test suite minimization approach, called ATM, which relies on test code similarity and evolutionary search. ATM preprocessed test code by removing the information that is irrelevant to the testing rationale and then converted it into Abstract Syntax Trees (ASTs). Then, four different tree-based similarity measures (i.e., top-down, bottom-up, combined, and tree edit distance) were used to calculate similarity values between these ASTs. Finally, evolutionary search (i.e., GA and NSGA-II) was employed using the similarity values as fitness to find, for a given test budget, an optimal subset of the test suite that contains diverse test cases. Results showed that the best configuration of ATM (i.e., GA with combined similarity) achieved a better trade-off in terms of effectiveness and efficiency of test suite minimization than FAST-R by achieving a significantly higher average fault detection rate ($+19\%$) while running within practically acceptable time ($1.2~{}hours$ on average), thus making it a better option in many contexts compared to FAST-R.
However, ATM suffers from limitations regarding its scalability for very large projects in the dataset, such as $Time$, which consists of $30k$ lines of code in its most recent version with large test suites (nearly $4k$ test cases per version), as minimization took more than $10$ hours per version as compared to FAST-R ($2.17$ seconds) and random minimization ($0.006$ seconds). The similarity measures used by ATM had a direct impact on its scalability. First, the similarity calculation time took up to $41.20\%$ (using tree edit distance) of the overall minimization time of ATM, which is due to the fact that converting test code into ASTs and calculating tree-based similarity based on ASTs are both time- and resource-consuming. Second, the search time, which was influenced by the employed similarity measures and the number of test cases per version, increased rapidly with the test suite size. Interestingly, though tree edit distance took much longer to calculate than other ATM similarity measures, it enabled GA search to converge faster towards the termination criterion (a fitness improvement of less than $0.0025$ across generations). Although GA with combined similarity was identified as the best ATM configuration based on the average minimization time across projects ($1.2~{}hours$), it is important to note that it was $2.5~{}hours$ slower than GA with tree edit distance for the largest test suite with nearly $4k$ test cases. Therefore, when dealing with large test suites, adopting a more informative similarity measure that facilitates faster search convergence can result in a notable reduction in minimization time, regardless of the computational cost associated with the similarity calculation.
In contrast to the above TSM approaches, LTM relies on test code without requiring any preprocessing, and employs large pre-trained language models (i.e., CodeBERT, GraphCodeBERT, and UniXcoder) and commonly used vector-based similarity measures (i.e., Cosine similarity and Euclidean distance) [33] to generate more informative similarity values between test case pairs. The goal is to both calculate similarity more efficiently and better guide the search algorithm (GA) to converge faster to a better fault detection capability.
3 LTM: Language Model-based Test Suite Minimization
This section describes our approach for test suite minimization, called LTM, relying on language models to compute test code similarity and enable the use of evolutionary search for minimization. Our goal is to find a black-box solution that achieves a better trade-off, in terms of scalability and effectiveness, than the latest state-of-the-art (SOTA) approach, namely ATM [8]. Despite the higher fault detection rates achieved by ATM, it has limitations in terms of scalability, which are due to the transformation of test cases into ASTs and the time-consuming calculation of tree-based similarity, making it difficult to apply for very large software systems. Similar to ATM, LTM is also a similarity-based approach operating under the assumption that there exists a positive correlation between test suite diversity and its fault detection capability [2]. We aim to find a similarity measure that is not only more computationally efficient but also provides better guidance for the search to converge faster to a higher fault detection capability and thus result in less search time. To achieve this, we employed pre-trained language models to generate code embeddings and used them as the basis to measure the vector-based similarity between test case pairs. The computation of vector-based similarity is much more efficient than tree-based similarity requiring tree traversal. Moreover, language models have the capability to capture patterns based on the syntax and semantics of test case code without requiring preprocessing or feature extraction, making them a more suitable alternative in our context.
Figure 1: Main steps of LTM to perform test suite minimization
Figure 1 outlines the main steps of LTM to minimize test suites. Regarding the first two steps, we describe how we tokenized the source code of test cases (Section 3.1.1) and then extracted embeddings using language models (Section 3.1.2). Then, we describe the algorithms (Cosine Similarity and Euclidean Distance) we employed for measuring the similarity between these embeddings (Section 3.2). Finally, we describe the search algorithm (Genetic Algorithm) we employed to minimize test suites using the similarity measures as fitness.
3.1 Language Models for Test Method Representation
ATM used ASTs (represented in XML format) to preserve the syntactic structure of test case code and calculated tree-based similarity between the ASTs of test case pairs, which we observed to be time- and resource-consuming due to the need for tree traversal, thus leading to scalability issues. In addition, we found that among all tree-based similarity measures, ATM using tree edit distance achieved the highest FDR and required the least search time across all project versions, since it made the search algorithm converge faster towards the termination criterion, that is fitness improvement is less than $0.0025$ across generations. This suggests that, regardless of the computational cost associated with its calculation, tree edit distance is a more informative similarity measure as it can be more effective in guiding the search to converge faster towards a higher fault detection rate. Therefore, we aim to develop a solution that overcomes the scalability issues of ATM by finding similarity measures that (1) do not require test code preprocessing, (2) are computationally efficient, and (3) are more informative, offering better guidance to the search algorithm and accelerating its convergence, thus reducing search time.
In this paper, we employ five pre-trained language models, namely CodeBERT [10], GraphCodeBERT [11], UniXcoder [12], StarEncoder [13], and CodeLlama [14]. We select these models for the following reasons:
  • CodeBERT, GraphCodeBERT, and UniXcoder are open-source publicly available language models that are specifically designed for programming languages and pre-trained on a large source code corpus. These language models surpassed the performance of other SOTA models in code clone detection tasks [10, 11, 12, 34], relying on the similarity between pairs of code fragments[12], thus making them viable options for similarity-based test suite minimization.
  • Similar to CodeBERT and its follow-up models, StarEncoder is a language model designed for understanding and analyzing code. Though it was not assessed for code clone detection tasks, StarEncoder is more recent and was pre-trained on a dataset encompassing more diverse ($86$) programming languages, enabling it to recognize test case patterns written in different coding styles.
  • Unlike the above models that are pre-trained for code understanding tasks, CodeLlama is specialized for code generation and infilling tasks (i.e., filling a missing part of a code snippet). It has $7Billion$ parameters, which is significantly higher than the parameter size ($125Million$) of the above models, and was pre-trained on a larger dataset, making it potentially able to distinguish test code more precisely.
In our context, these language models take the source code of test cases as input, without requiring any preprocessing or transformation into other formats, and generate embeddings that capture both semantic and contextual information from test case code. The output of these language models is represented as numeric vectors, which offer the opportunity to employ a variety of vector-based algorithms to measure similarities between test cases. The process starts by passing the source code of a test case (a Java test method in our context) as input to the language models, which is then converted to a list of tokens, each of which is assigned an individual numerical vector representation. The final output is a vector representation (embedding) generated for the test method, which aggregates the information from all individual vector representations, which is called a test method embedding in our context.
3.1.1 Tokenizing Test Methods
Language models deal with a test code fragment (method) as a sequence of tokens and split it using byte-pair-encoding (BPE) [35], which is a subword segmentation algorithm that splits words into a sequence of sub-words. This enables language models to better handle out-of-vocabulary words, such as method and variable names. For example, the tokenizer splits the method name testCloning into Ġtest (where Ġ denotes a space), Cl and oning, since the testCloning word does not exist in the vocabulary and is thus separated into these three sub-words. The generated tokens are then processed as follows.
  • CodeBERT, GraphCodeBERT, and StarEncoder add two special tokens, namely $[CLS]$ and $[SEP]$, to the beginning and end of each sequence of tokens, respectively. The $CLS$ token is a special token that represents the whole input sequence, whereas the $[SEP]$ token is a separator token that denotes the end of the sequence [23]. In summary, each test method is represented as $[CLS],c_{1},c_{2},...,c_{m},[SEP]$, where $c_{i}$ denotes the $i^{th}$ code token and $m$ is the total number of code tokens.
  • For UniXcoder, besides the $[CLS]$ and $[SEP]$ tokens, an additional special token ($[Enc]$, $[Dec]$, or $[E2D]$) is added to the beginning of the input indicating the pre-training mode of the model as encoder-only, decoder-only, or encoder-decoder mode, respectively. These modes differ in terms of model architectures and training tasks during pre-training, thus supporting various downstream tasks. We used the encoder-only mode for producing contextualized code embeddings, as decoder-related modes are used for code generation [12]. Therefore, the final input for UniXcoder is represented as $[CLS],[Enc],[SEP],c_{1},c_{2},...,c_{m},[SEP]$.
  • CodeLlama uses $<$$s$$>$ and $<$$/s$$>$ tokens denoting the beginning and end of each token sequence, respectively: ($<$$s$$>$$,c_{1},c_{2},...,c_{m},$$<$$/s$$>$).
During pre-training, each token is then mapped to a $768$-dimensional vector representation that contains the semantic and contextual information of this token for all language models above, except for CodeLlama where the vector size is $4,096$. We set the token length to $512$ for all language models during the tokenization process.
3.1.2 Generating Test Method Embeddings
A test method embedding is a numerical vector representation of a test method that captures semantic and contextual information from the source code. It is based on how and on what data a model was pre-trained, as follows.
  • CodeBERT pre-trained code representations are based on a large corpus called CodeSearchNet [36], which contains both natural language (e.g., code comments) and source code across six programming languages. Its follow-up models (GraphCodeBERT and UniXcoder) leveraged data flow information during pre-training, which captures relationships between variables in the input code fragments and ASTs of the source code, respectively, to enhance the code representation.
  • StarEncoder was pre-trained on a large dataset encompassing 86 programming languages collected from The Stack [37], which is specifically crawled from GitHub repositories for pre-training code LLMs.
  • CodeLlama was pre-trained on a massive dataset ($500Billion$ tokens) containing both source code and natural language. CodeLlama has multiple variations and model sizes, depending on the type of training data and parameter size. In this paper, considering the importance of scalability, we used the smallest version of CodeLlama (i.e., CodeLlama-7b), which has seven billion parameters.
In terms of model architecture:
  • CodeBERT and StarEncoder employ a multi-layer bidirectional self-attentive Transformer [38] as model architecture to help the model capture contextual and positional information for each token from the entire input sequence.
  • GraphCodeBERT extend such architecture using a graph-guided masked attention function to help the model encode graph-based data flow information.
  • UniXcoder, on the other hand, is based on a multi-layer Transformer with a prefix denoting the pre-training mode of the model. The model architecture for the encoder-only mode, which is employed by LTM, allows the model to learn, for each token, the contextual information from the entire input sequence.
  • CodeLlama uses an optimized auto-regressive transformer, enabling the model to efficiently predict missing parts within code sequences. Despite being a decoder-only model, CodeLlama exhibits the capability to understand the test code context and generate missing parts accurately.
During pre-training:
  • CodeBERT employs Masked Language Modeling (MLM) [23], which allows the model to predict the masked tokens from the input sequence and whether a token was randomly replaced in a given sequence [39]. These two tasks help the model generate code embeddings that contain more accurate contextual information, accounting for the position of the tokens in the source code input.
  • GraphCodeBERT also employs MLM with two additional tasks: Edge Prediction [11] and Node Alignment [11]. Edge Prediction allows the model to predict which edges, which denote dependencies between variables, have been masked for a given variable in the data flow graph. Node Alignment allows the model to predict which code token is related to a given variable in the data flow graph. These two tasks further enhance the code representation using data flow information.
  • UniXcoder also uses the MLM task for the encoder-only mode, and further learns the test method embedding, which is first generated by taking the average (i.e., mean pooling) of all token embeddings and then further trained by utilizing multi-modal contrastive learning [40] and cross-modal generation [12], which both leverage AST and code comments to further enhance test method embeddings. Multi-modal contrastive learning generates different embeddings for the same input but with different dropout masks, which allows the model to predict the original embedding using the other generated embeddings. This task helps the model to better distinguish between different method embeddings and thus can better deal with downstream tasks, such as test method similarity analysis [40]. Cross-modal generation generates code comments for the input test method. This helps the model to fuse the semantic information from the code comments, which describe the function of the code, to the test method embeddings.
  • StarEncoder also employs MLM in addition to Next Sentence Prediction (NSP) [23], which helps the model learn the relationship between test code segments by predicting whether segments follow other segments.
  • Unlike the above models, CodeLlama employs code infilling task, which allows the model to generate the missing part of a code snippet while comprehending the entire context. This enables the model to understand the logical structure of the test code and the relationships between test code segments.
For LTM using CodeBERT, GraphCodeBERT, and StarEncoder, we use the embedding that corresponds to the $[CLS]$ token as a test method embedding, since it aggregates the information from all tokens of that method. For LTM using UniXcoder, we use the pre-trained test method embedding corresponding to each test method. For LTM using CodeLlama, we rely on the mean pooling of the last hidden states (i.e., the average of all token embeddings extracted from the last hidden layer of the model), which is also an effective strategy that aggregates the information from all code tokens [41].
3.2 Similarity Measurement of Test Method Embeddings
LTM employs two similarity measures for calculating the similarity between test method embeddings: Cosine Similarity and Euclidean Distance. They measure the similarity between test method embeddings from different aspects: Cosine similarity measures the angle between two vectors, whereas Euclidean distance calculates the straight-line distance between them.
Cosine similarity. This is a measure of similarity between two vectors based on the cosine of the angle between them [42], which is the dot product of the vectors divided by the product of their lengths, and is calculated as follows:
$Cosine~{}Similarity=\frac{\mathbf{T_{1}}\cdot\mathbf{T_{2}}}{\|\mathbf{T_{1}}\|\|\mathbf{T_{2}}\|}$ (1)
where $\mathbf{T_{1}}$ and $\mathbf{T_{2}}$ denote the embeddings of test case $T_{1}$ and $T_{2}$, respectively.
The value of Cosine similarity ranges from $-1$ to $1$. The higher the value of Cosine similarity, the more similar the two test cases are. In order to bound the value of Cosine similarity between 00 and $1$, we normalized it as follows:
$Norm.~{}Cosine~{}Sim.=1-\frac{\arccos{\frac{\mathbf{T_{1}}\cdot\mathbf{T_{2}}}{\|\mathbf{T_{1}}\|\|\mathbf{T_{2}}\|}}}{\pi}$ (2)
Euclidean distance. This is a measure of similarity between two vectors based on the square root of the sum of squared differences between corresponding elements of the two vectors [42], which is calculated as follows:
$Euclidean~{}Distance=(\sum_{i}^{m}{(t_{1i}-t_{2_{i}})^{2}})^{1/2}$ (3)
where $t_{1i}$ and $t_{2i}$ is the $i^{th}$ element of embedding $T_{1}$ and $T_{2}$, respectively. $m$ is the total number of elements in each embedding ($4,096$ for CodeLlama and $768$ for the other models).
The value of Euclidean Distance ranges from 00 to $\infty$. The higher the value of Euclidean distance, the less similar the two test cases. In order to bound the Euclidean distance between 00 and $1$, and obtain a similarity measure, we normalized the distance as follows [43]:
$Norm.~{}Euclidean~{}Dist.=\frac{1}{1+(\sum_{i}^{m}{(t_{1i}-t_{2_{i}})^{2}})^{1/2}}$ (4)
We used the ‘pdist’ function [^1] https://docs.scipy.org/doc/scipy/reference/generated/scipy.spatial.distance.pdist.html from the Scipy Python library, which performs pairwise calculations for large datasets. The ‘pdist’ employs highly optimized C code to improve the efficiency of similarity calculation on vector-based data, and stores the output in a condensed matrix, which further reduces the required memory resources.
3.3 Search-based Test Suite Minimization
Given that test suite minimization is an NP-hard problem [2], it can be addressed efficiently using meta-heuristic search algorithms [8] to find feasible, near-optimal solutions, a minimized test suite containing diverse test cases for a given budget in our context. Like ATM, we relied on a Genetic Algorithm (GA) since it has shown, in the context of TSM, to achieve a better trade-off between effectiveness and efficiency than its multi-objective alternative, namely Non-Dominated Sorting Genetic Algorithm II (NSGA-II) [8]. The optimization problem addressed by the GA is defined as a fixed-size subset selection problem [44]. Each solution (chromosome) is a subset and is represented as a binary vector where $1$ denotes the selection of the test case and 00 otherwise. The vector length equals the total number of test cases in the test suite before minimization. Given specified minimization budgets (25%, 50%, and 75%), the percentage of selected test cases in each solution is set to equal to the budget. For the fitness function, we used the summation of the maximum squared similarity values (i.e., Normalized Cosine Similarity and Normalized Euclidean Distance), shown below:
$Fitness=\frac{\sum_{i,t_{i}\in M_{n}}(Max_{i,j,t_{i},t_{j}\in M_{n},i\neq j}\ Sim(t_{i},t_{j}))^{2}}{n}$ (5)
where $M_{n}$ is a minimized test suite of size $n$, and $Sim(t_{i},t_{j})$ is the normalized similarity score of test cases pair $t_{i}$ and $t_{j}$.
To select hyperparameter values, we followed the same published guidelines [^2] https://www.obitko.com/tutorials/genetic-algorithms/recommendations.php as ATM for GA search, by using a population size of $100$, a mutation rate of $0.01$, and a crossover rate of $0.90$. The GA process maximizes the diversity of the subset by iteratively evolving the population using crossover and mutation operators while evaluating each subset using the predefined fitness function. For each generation, the size of each solution (subset) remains fixed and equal to the minimization budget. This process is repeated until the fitness value improvement across generations is less than $0.0025$.
Figure 2: An example of how similarity values of pairs of test cases are represented using a matrix. Values on and below the diagonal are set to 0 as they are either useless or duplicates.
In terms of implementation, LTM optimizes GA search by changing the data structure of the input (i.e., similarity values) to (1) accelerate the fitness calculations and (2) enhance memory usage. The data structure employed by ATM for storing similarity values is a data frame  [^3] https://pandas.pydata.org/docs/reference/api/pandas.DataFrame.html, which has a complex indexing and labeling mechanism, thus making fitness calculation computationally expensive. Therefore, we use matrices  [^4] https://numpy.org/doc/stable/reference/generated/numpy.array.html instead, since they facilitate faster numerical calculations due to the fact that (1) it is a simpler data structure without complex indexing and labeling and (2) the matrix operations are implemented using highly optimized C code that ensures efficient computation and memory management. As shown in Figure 2, we stored the similarity values for all pairs of test cases in a matrix format without any duplicate and useless values (the similarity values on and below the diagonal of the matrix were set to 0). To save memory and further accelerate calculations, we employed a sparse matrix format  [^5] https://docs.scipy.org/doc/scipy/reference/sparse.html, which exclusively preserves the value and position (indicating for which pair of test cases a given similarity value represents) of non-zero similarity values. Note that we did not parallelize the fitness calculation as it significantly increases the memory required by the matrix, which must contain similarity values for all considered subsets, making it less scalable. Although fitness calculations can be parallelized in other ways, such as multiprocessing, which utilizes multiple CPUs to parallelize fitness calculation, or distributed computing [44], which distributes the fitness calculation to various devices or nodes, they require additional computational resources, which contradicts our goal of improving scalability.
4 Validation
This section reports on the experiments we conducted to assess the effectiveness and efficiency of LTM.
4.1 Research Questions
RQ1: How does LTM perform for test suite minimization under different configurations? Test suite minimization aims to reduce the number of test cases in a test suite by removing redundant test cases while maintaining its fault detection power, thus achieving higher testing efficiency and effectiveness. In particular, the performance of LTM can be influenced by (a) the language models used for extracting test method embeddings and (b) the distance functions used for measuring the similarity between test method embeddings. In this RQ, we assess the performance, in terms of effectiveness and efficiency, of LTM under different configurations, each with a different combination of language model and distance function. We evaluated LTM using three minimization budgets (i.e., $25\%$, $50\%$, and $75\%$).
  • RQ1.1: How effectively can LTM minimize test suites? The language models used for LTM (i.e., CodeBERT, GraphCodeBERT, and UniXcoder) vary in terms of architectures and training data, thus generating embeddings that capture different information about test case code. The distance functions we used to measure the similarity or distance between test method embeddings differently, which can lead to different comparison results. Therefore, the combination of language models and distance functions can affect the ability of LTM to remove redundant test cases, which can in turn affect its fault detection capability. In this RQ, we assess the effectiveness of LTM in terms of fault detection capability under different configurations. In addition, we calculated the execution time of minimized test suites to quantify the achieved time savings resulting from test suite minimization.
  • RQ1.2: How efficiently can LTM minimize test suites? Test suite minimization should run within practical time. The use of different combinations of language models and distance functions can impact the preparation time, and the resulting similarity values can guide the GA search to varying degrees of efficiency, thus resulting in different search time. In this RQ, we assess the efficiency of LTM in terms of preparation time (taken for test code tokenization, embeddings extraction, and similarity calculation) and search time (taken for GA search).
RQ2: How does LTM compare to ATM? ATM was developed as a black-box test suite minimization approach that achieved a better trade-off, in terms of effectiveness and efficiency, than the SOTA approach (FAST-R [6]). However, the minimization time of ATM increased rapidly with test suite size, making it less scalable for large projects with large test suites. For example, ATM required more than $10$ hours, in terms of total minimization time, to minimize a large project in the dataset used for its evaluation, thus indicating scalability issues. In this RQ, we assess the performance of LTM compared to the two best configurations of ATM, specifically ATM using GA with combined and tree edit distance similarity measures, in terms of both effectiveness and efficiency.
  • RQ2.1: How effectively can LTM minimize test suites compared to ATM? We aim to find an informative similarity measure that can better capture similarities between test cases, thus achieving higher FDR. In this RQ, we compare the best configuration of LTM to the two best configurations of ATM in terms of fault detection capability.
  • RQ2.2: How efficiently can LTM minimize test suites compared to ATM? To achieve better scalability compared to ATM, we aim to reduce (a) preparation time, by investigating a similarity measure that is more efficient to calculate, and (b) search time, by investigating a more informative similarity measure that can better guide the search, thus resulting in faster convergence. In this RQ, we compare the efficiency of LTM to the two best configurations of ATM in terms of both preparation time and search time. Further, to better assess scalability for both LTM and ATM, we fit a regression model to investigate how minimization time (the sum of preparation time and search time) increases with the size of a test suite.
4.2 Experimental Design and Dataset
We conducted experiments to evaluate the performance of alternative LTM configurations, each with a different combination of language model and similarity measure, resulting in ten configurations in total (five different language models and two similarity measures), using the same experimental design and dataset as ATM. We conducted our experiments on a cluster of $1,340$ nodes with $83,216$ available CPU cores, each with a $2x$ AMD Rome 7532 with 2.40 GHz CPU, 256M cache L3, 249GB RAM, running CentOS 7. We considered each Java project version as an independent subject and conducted our experiments on all Java project versions ($835$ in total) in the dataset to increase the number of instances for experimental evaluation and, therefore, enhance the soundness of our conclusions. Each TSM approach was performed on every project version and ran $10$ times to account for randomness and draw statistically valid conclusions, using three minimization budgets ($25\%$, $50\%$, and $75\%$), which is the percentage of test cases that retained in the minimized test suite). There is a total of $25,050$ runs ($835$ Java project versions $\times$ $10$ runs per version $\times$ $3$ minimization budgets) for each configuration. Overall, the experiments, including similarity calculations, GA search, as well as executing all test suites $10$ times, took more than three months in calendar time corresponding to 6 years of computation. We identified the best configuration of LTM considering both the effectiveness and efficiency and compared it to the two best configurations of ATM.
4.2.1 Baseline approach (ATM)
We compared LTM to ATM [8], a similarity-based, search-based test suite minimization approach that relies solely on the source code of test cases. ATM preprocesses test code by removing information that is irrelevant to testing logic from the test code, then converts test code into ASTs. ATM calculates four different tree-based similarity measurements between test case ASTs: top-down, bottom-up, combined and tree edit distance similarity. Search algorithms (i.e., GA and NSGA-II) were then employed based on similarity values to find an optimal subset of test cases by removing redundant test cases. Using GA with combined similarity, ATM achieved a better trade-off in terms of effectiveness and efficiency compared to a SOTA test suite minimization approach, FAST-R [6]. However, it suffered from scalability issues as, for example, it took more than $10$ hours to minimize $4k$ test cases for the largest test suite. This motivated us to compare LTM to ATM, in terms of fault detection capability and, more importantly, minimization time. We aim to assess whether LTM can be more scalable than ATM while achieving comparable or even higher fault detection capability. We compared LTM to the two best configurations of ATM, using GA with combined similarity (ATM/Combined) and tree edit distance (ATM/TreeEditDistance). The former was the best configuration considering the trade-off between effectiveness and efficiency, whereas the latter achieved higher fault detection capability but was more time-consuming than the former in terms of total minimization time. We further evaluated the achieved saving in testing time for every minimized test suite for both LTM and ATM.We relied on the publicly available replication package of ATM [^6] https://doi.org/10.5281/zenodo.7455766 provided by its authors.
4.2.2 Dataset
To facilitate comparisons, we ran our experiments using the same dataset as ATM to evaluate the effectiveness and, more importantly, the scalability of LTM. The dataset consists of $16$ Java projects collected from Defects4J, a well-known dataset that offers real and reproducible faults extracted from real-world, open-source Java projects to support software testing research. There is no publicly available industrial system or open source system that contains information to automatically trace test failures to system faults, a key requirement for our experiments. To further validate the performance of LTM, we included an additional large project from Defects4J, called $Closure$, which was not part of the dataset used to evaluate ATM due to scalability issues. Each project has numerous faulty versions, each of which contains a single real fault that is present in the production code and leads to the failures of one or more test cases. It is worth noting that there is currently no publicly available dataset that offers multiple real faults per version, as automatically establishing a clear link between faults and test case failures would be challenging. TableI presents the characteristics of all $17$ projects. Overall, the project sizes in terms of KLoc range from $2$ to $179$ KLoc, with the corresponding tests sizes ranging from $4$ to $253$ KLoc. The number of faulty versions ranges from $4$ to $174$ across projects, with the average number of test cases per version ranging from $152$ to $7,308$ across projects. Project sizes and tests sizes (i.e., lines of code) were extracted by analyzing the most recent version of each project using the CLOC [^7] https://github.com/AlDanial/cloc tool. In summary, we evaluated LTM using a total of $835$ versions from $17$ projects. However, given the scalability limitations of ATM [8], we compared our results to ATM using only the $661$ versions from the original $16$ projects, excluding the $Closure$ project.
TABLE I: Summary of the $16$ Java projects
Project Project Size # of versions Average # of test cases Tests Size
(KLoC) (faults) (methods) per version (KLoC)
Chart 92 26 1,817 41
Cli 2 39 256 4
Codec 9 18 413 15
Collections 30 4 1,040 38
Compress 45 47 404 29
Csv 2 16 193 7
Gson 9 18 984 20
JacksonCore 31 26 356 45
JacksonDatabind 74 112 1,814 72
JacksonXml 6 6 152 10
Jsoup 14 93 494 13
JxPath 20 22 250 6
Lang 30 64 1,796 61
Math 71 106 2,078 73
Mockito 21 38 1,182 36
Time 30 26 3,918 56
Closure 179 174 7,308 253
4.2.3 Evaluation metrics
We used the same evaluation metrics as ATM, specifically fault detection rate and total minimization time, in addition to the achieved saving in testing time of minimized test suites.
Fault Detection Rate (FDR). We use FDR to assess the effectiveness of LTM compared to ATM. Given that each project version contains a single fault, the corresponding fault detection rate per version can either be $1$ (fault detected) or 00 (fault undetected). Therefore, we calculate the fault detection rate for each project by considering all its versions, as follows:
$FDR=\frac{\sum_{i=1}^{m}f_{i}}{m}$ (6)
where $m$ denotes the total number of versions of the project. For a project version $i$, $f_{i}$ is equal to $1$ if at least one failing test case is included in the minimized test suite, indicating the detection of the fault of that version, or 00 otherwise.
For example, in the $Chart$ project, which has a total of $26$ faulty versions, if the minimized test suites detected the faults in 21 versions but did not in $5$ versions, FDR would be calculated as $21/26$, resulting in an FDR of $0.81$. Note that we ran our experiments for each version a total of $10$ times to mitigate randomness. Hence, the final FDR for each project is obtained by taking the average FDR results across the $10$ runs.
Total Minimization Time (MT). We assess the efficiency of LTM compared to ATM by computing (1) the preparation time, which includes the time required for loading the language models, tokenizing test case code, extracting test method embeddings, and calculating similarity values between all pairs of test cases, and (2) the search time, which is the time required for running the GA.
Time Saving Rate (TSR). We use TSR to quantify the achieved testing time savings resulting from test suite minimization. It is important to note that the reduction in test execution time may not be directly proportional to the reduction in the test suite size, as the time required to execute individual test methods may vary [1]. Using the Defects4J infrastructure, we ran the test suite for each project version ($835$ versions in total) $10$ times and collected the average execution time for each test method in a clean, stable environment, a cluster with many available CPU cores. This ensures that each run runs independently and is not impacted by external interferences and fluctuations in memory or CPU performance. Then, we calculate the TSR for each minimized test suite as follows:
$TSR=(1-\frac{test~{}execution~{}time~{}after~{}TSM}{test~{}execution~{}time~{}before~{}TSM})*100$ (7)
Fisher’s exact test. We use Fisher’s exact test [45], a non-parametric statistical hypothesis test, to assess whether the difference in proportions of detected faults between the alternative LTM and ATM configurations is significant.
4.3 Results
In this section, given space constraints, we only report results for a minimization budget of 50% (the percentage of test cases retained in the minimized test suite) as trends and conclusions are similar for other budgets (25% and 75%).
4.3.1 A preliminary analysis of using full test code versus preprocessed test code
One important question is whether the source code of test cases requires preprocessing, such as removing test case names, logging statements, comments, and test assertions. Though such preprocessing was performed by ATM [8], this information may still convey relevant information about test case similarity, thus retaining them can improve fault detection capability. This information can also help language models better learn about semantic and contextual information about test cases and thus generates more informative test method embeddings. For example, code comment ‘Tests the equals method’ describes the functionality of the test code, thus providing potentially important semantic information [12]. Therefore, we conducted a preliminary study to compare the FDR results using full code (i.e., the source code of test cases with no preprocessing) and preprocessed code (as done by ATM) on a sample of $10$ projects from the Defects4J dataset. We observed that LTM achieved higher FDR results using full code than using prepossessed code ($+0.06$ on average) across all configurations, thus making it a better choice.
Note that we did not consider using the full code for ATM. Using unprocessed code would make the preparation phase of ATM much more time- and resource-consuming. This is due to the fact that (1) the xml files used to store Abstract Syntax Trees (ASTs) of test cases require substantial storage space and (2) it would take longer time to traverse the ASTs of test cases, thus yielding longer preparation time for transforming the test case code into ASTs and calculating similarity values. Given that, for ATM, preparation time can take up to 41.2% of the total minimization time, using the full code for ATM would make matters much worse.
4.3.2 Improvement using optimized GA.
As discussed in Section 3.3, we optimized GA to accelerate the search process. Our results revealed that, compared to MT before optimization, the search time of ATM/Combined and ATM/TreeEditDistance is $224.66$ and $179.30$ times faster per project version, respectively. For LTM with UniXcoder/Cosine, the search time is $272.90$ times faster per project version, a significant improvement in scalability. Note that the following comparisons between LTM and ATM are based on applying this optimized GA to both approaches.
4.3.3 RQ1 results
Tables II, III, and IV provide the results of LTM in terms of FDR, MT (in min), and TSR (in percentage), respectively, when using CodeBERT, GraphCodeBERT, UniXcoder, StarEncoder, and CodeLlama-7b with Cosine and Euclidean similarity measures, for a $50\%$ minimization budget.
RQ1.1 results. Table II shows that all LTM configurations achieved a high FDR, ranging from $0.65$ to $0.95$ across projects, with both the mean and median ranging from $0.75$ to $0.84$ for all configurations, for a $50\%$ minimization budget. The highest overall average FDR was achieved using UniXcoder/Cosine (mean = $0.84$, median = $0.82$), ranging from $0.75$ to $0.95$ across projects. However, LTM using CodeBERT/Euclidean also yielded a high FDR (mean = $0.82$, median = $0.84$). CodeBERT/Euclidean and UniXcoder/Cosine showed standard deviations of $0.06$ and $0.07$, respectively. In addition, Fisher’s exact test results reveal that LTM using UniXcoder/Cosine achieved significantly better FDR results than all other configurations, with $\alpha=0.01$.
Figure 3 depicts the average FDR for all projects across generations, for all configurations of LTM. Compared to other configurations, LTM using UniXcoder/Cosine converges faster to a higher FDR ($0.84$ across projects), suggesting that it offers better guidance for the GA search, making it more effective at removing redundant test cases and thus leading to a better fault detection rate.
We observed that, for each language model, there seems to be a significant difference in FDR between using Cosine Similarity and Euclidean Distance, which is further confirmed by Fisher’s exact test with $\alpha=0.01$. This suggests that FDR results are influenced by the combination of language models and distance functions. The embeddings generated by UniXcoder and GraphCodeBERT are enriched with detailed information about the nodes and edges within test code ASTs. This could be the reason that Cosine similarity, as a directional similarity, is more informative in distinguishing such test code embeddings than Euclidean Distance, a magnitude similarity. For CodeLlama-7b, the embedding dimensionality is $4,096$, compared to $768$ for other language models, which can explain its better performance with Cosine Similarity over Euclidean Distance, since the latter is less effective in high-dimensional spaces [46].
In certain projects ($Cli$, $Codec$, $Csv$, and $Time$), UniXcoder/Cosine yielded a $5\%$ to $10\%$ lower FDR than other configurations. However, Fisher’s Exact Test revealed that those differences are not significant. Nevertheless, we provide some tentative explanations regarding these projects: (1) A larger training data set covering a greater variety of programming languages (StarEncoder) and a substantially greater number of parameters (CodeLlama) might be beneficial for understanding test code from these projects, and (2) using ASTs as input for pre-training might introduce unnecessary details into test code embeddings, if test cases in these projects are similar in terms of code structure but rather differ regarding low-level details (e.g., complex string manipulation or domain-specific character strings), thus making UniXcoder less effective in understanding their test code.
In addition, we observed that UniXcoder/Cosine achieved higher average FDR than CodeLlama-7b/Cosine (mean = $0.80$, median = $0.81$) and CodeLlama-7b/Euclidean (mean = $0.76$, median = $0.76$). This suggests that language models with fewer parameters ($125Million$ for UniXcoder) can outperform much larger language models ($7Billion$ for CodeLlama-7b) for our minimization problem. This can be attributed to the fact that CodeLlama-7b, like most large language models, is a decoder-only model that was pre-trained for code generation tasks, while UniXcoder employs code understanding tasks that are more effective in analyzing and comprehending the contextual semantic information from the source code [47].
Recall that the GA terminates when the fitness value improves by less than $0.0025$, thus resulting in different numbers of generations for each configuration across project versions. The overall average number of generations of all configurations across versions is $47.73$, with UniXcoder/Cosine having a higher average ($63.10$ generations). However, if a fixed number of generations had been used, say $40$ generations, UniXcoder/Cosine would still yield the highest FDR ($0.81$) among all configurations. This clearly makes UniXcoder/Cosine the best LTM alternative, in terms of FDR.
Figure 3: FDR across projects for each generation of LTM and ATM
TABLE II: Results and descriptive statistics of FDR of LTM and ATM across projects for the $50\%$ minimization budget. The highest FDR results are highlighted in bold.
 Project Approach CodeBERT GraphCodeBERT UniXcoder StarEncoder CodeLlama-7b ATM ATM
Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Euclidean Combined Tree Edit Distance
Chart 0.82 0.84 0.79 0.75 0.82 0.69 0.71 0.79 0.72 0.73 0.88 0.89
Cli 0.81 0.81 0.83 0.86 0.82 0.81 0.83 0.86 0.87 0.86 0.87 0.87
Codec 0.89 0.90 0.83 0.88 0.80 0.81 0.82 0.83 0.87 0.78 0.77 0.80
Collections 0.70 0.82 0.95 0.75 0.95 0.70 0.70 0.75 0.78 0.70 0.98 0.92
Compress 0.85 0.86 0.87 0.87 0.86 0.81 0.78 0.84 0.85 0.79 0.88 0.89
Csv 0.86 0.87 0.88 0.91 0.85 0.79 0.84 0.86 0.93 0.88 0.92 0.91
Gson 0.70 0.71 0.72 0.75 0.81 0.75 0.76 0.76 0.81 0.77 0.76 0.77
JacksonCore 0.70 0.73 0.73 0.73 0.75 0.67 0.74 0.74 0.75 0.75 0.69 0.72
JacksonDatabind 0.66 0.69 0.68 0.67 0.76 0.65 0.65 0.69 0.67 0.65 0.69 0.70
JacksonXml 0.77 0.85 0.83 0.73 0.83 0.82 0.67 0.65 0.83 0.85 0.60 0.73
Jsoup 0.75 0.74 0.77 0.76 0.80 0.76 0.73 0.75 0.77 0.75 0.70 0.69
JxPath 0.88 0.90 0.92 0.86 0.94 0.83 0.72 0.77 0.78 0.75 0.79 0.79
Lang 0.81 0.85 0.86 0.75 0.90 0.71 0.77 0.85 0.83 0.71 0.78 0.83
Math 0.74 0.78 0.75 0.74 0.78 0.65 0.71 0.78 0.75 0.67 0.81 0.82
Mockito 0.80 0.83 0.84 0.81 0.91 0.79 0.77 0.83 0.80 0.77 0.77 0.79
Time 0.83 0.88 0.85 0.82 0.82 0.77 0.82 0.79 0.81 0.76 0.89 0.88
             Statistics
Min 0.66 0.69 0.68 0.67 0.75 0.65 0.65 0.65 0.67 0.65 0.60 0.69
25% Quantile 0.73 0.77 0.77 0.75 0.80 0.70 0.71 0.75 0.77 0.73 0.75 0.76
Mean 0.79 0.82 0.82 0.79 0.84 0.75 0.75 0.78 0.80 0.76 0.80 0.81
Median 0.81 0.84 0.83 0.76 0.82 0.77 0.75 0.79 0.81 0.76 0.79 0.81
75% Quantile 0.84 0.86 0.86 0.86 0.87 0.81 0.79 0.83 0.84 0.78 0.88 0.88
Max 0.89 0.90 0.95 0.91 0.95 0.83 0.84 0.86 0.93 0.88 0.98 0.92
TABLE III: Results and descriptive statistics of MT (in min) of LTM and ATM across projects for the $50\%$ minimization budget. The shortest MT results are highlighted in bold.
 Project Approach CodeBERT GraphCodeBERT UniXcoder StarEncoder CodeLlama-7b ATM ATM
Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Euclidean Combined Tree Edit Distance
Chart 0.92 1.51 0.97 1.20 1.33 0.99 2.24 2.68 21.02 20.92 7.86 94.22
Cli 0.11 0.16 0.14 0.15 0.17 0.14 0.16 0.22 4.39 4.38 0.23 1.20
Codec 0.15 0.21 0.18 0.17 0.22 0.18 0.25 0.33 5.75 5.65 0.42 3.02
Collections 0.37 0.51 0.43 0.50 0.58 0.38 0.79 1.11 8.96 8.80 2.16 14.89
Compress 0.16 0.20 0.19 0.18 0.22 0.17 0.25 0.34 9.26 9.18 0.55 6.55
Csv 0.10 0.13 0.12 0.11 0.15 0.11 0.14 0.16 4.19 4.15 0.16 0.35
Gson 0.31 0.39 0.31 0.33 0.47 0.32 0.68 0.79 8.31 8.26 1.48 6.09
JacksonCore 0.14 0.16 0.16 0.15 0.19 0.15 0.22 0.25 11.10 11.06 0.35 2.13
JacksonDatabind 0.92 1.40 0.91 1.35 1.43 1.33 2.15 2.46 34.67 35.41 4.90 22.69
JacksonXml 0.11 0.11 0.11 0.10 0.13 0.11 0.12 0.13 6.86 6.84 0.13 0.34
Jsoup 0.17 0.17 0.17 0.17 0.22 0.18 0.29 0.29 20.92 20.91 0.55 1.63
JxPath 0.11 0.15 0.13 0.13 0.17 0.13 0.17 0.21 4.98 4.95 0.22 0.73
Lang 0.91 1.47 0.91 1.08 1.38 0.97 2.35 2.82 15.09 14.98 5.44 33.47
Math 1.33 2.33 1.31 2.06 2.12 1.98 3.41 4.22 21.12 23.09 12.05 215.43
Mockito 0.40 0.50 0.42 0.40 0.59 0.41 0.97 1.12 17.40 17.93 2.32 9.34
Time 3.49 5.35 3.43 3.91 3.73 3.44 9.06 10.45 41.75 39.60 26.13 163.18
             Statistics
Min 0.10 0.11 0.11 0.10 0.13 0.11 0.12 0.13 4.19 4.15 0.13 0.34
25% Quantile 0.13 0.16 0.16 0.15 0.19 0.15 0.21 0.24 6.58 6.54 0.32 1.52
Mean 0.61 0.92 0.68 0.75 0.82 0.69 1.45 1.72 14.74 14.72 4.06 35.95
Median 0.24 0.30 0.25 0.26 0.35 0.25 0.49 0.57 10.18 10.12 1.02 6.32
75% Quantile 0.91 1.42 0.91 1.11 1.34 0.98 2.17 2.52 20.95 20.91 5.04 25.39
Max 3.49 5.35 3.43 3.91 3.73 3.44 9.06 10.45 41.75 39.60 26.13 215.43
TABLE IV: Results and descriptive statistics of TSR (in percentage) of LTM and ATM across projects for the $50\%$ minimization budget. The highest TSR results are highlighted in bold.
 Project Approach CodeBERT GraphCodeBERT UniXcoder StarEncoder CodeLlama-7b ATM ATM
Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Euclidean Combined Tree Edit Distance
Chart 55.45% 57.35% 58.78% 59.81% 58.03% 56.11% 58.47% 63.08% 56.64% 59.14% 59.77% 57.53%
Cli 29.90% 27.90% 30.54% 29.75% 28.32% 28.17% 29.52% 27.52% 28.25% 28.46% 28.07% 27.34%
Codec 44.11% 40.68% 47.59% 49.93% 55.58% 50.62% 51.05% 48.38% 45.27% 48.83% 34.62% 45.94%
Collections 57.40% 55.05% 56.21% 56.10% 55.41% 59.05% 59.41% 57.36% 55.01% 56.78% 53.57% 57.65%
Compress 37.74% 37.98% 36.56% 38.10% 39.10% 40.70% 37.87% 35.26% 36.45% 40.11% 33.85% 32.59%
Csv 18.66% 18.78% 19.90% 20.07% 18.13% 10.11% 16.40% 16.15% 12.82% 9.37% 20.45% 19.89%
Gson 41.81% 40.60% 42.16% 42.01% 41.40% 44.63% 41.85% 40.10% 41.69% 43.63% 40.35% 38.17%
JacksonCore 47.45% 48.44% 44.37% 46.18% 47.10% 48.00% 46.02% 46.42% 46.95% 46.98% 43.59% 46.71%
JacksonDatabind 42.54% 41.28% 41.68% 41.63% 38.32% 42.37% 41.71% 39.56% 41.29% 41.75% 50.61% 47.92%
JacksonXml 42.81% 42.05% 42.63% 44.12% 45.80% 46.32% 45.70% 49.52% 44.03% 43.60% 39.17% 39.34%
Jsoup 42.96% 42.03% 41.33% 43.80% 33.90% 41.65% 38.60% 35.61% 35.12% 37.55% 40.39% 38.20%
JxPath 43.37% 42.82% 41.00% 41.90% 40.83% 42.70% 42.46% 41.84% 44.12% 44.16% 44.40% 44.46%
Lang 38.36% 36.49% 40.14% 41.24% 34.05% 44.76% 41.04% 35.37% 42.83% 44.97% 37.71% 38.06%
Math 50.24% 49.43% 50.26% 49.47% 49.67% 49.67% 50.00% 49.72% 50.90% 51.34% 47.93% 49.10%
Mockito 27.65% 29.87% 34.33% 38.38% 33.97% 39.93% 44.82% 41.80% 32.13% 39.77% 24.31% 26.13%
Time 48.12% 47.57% 48.33% 48.49% 47.89% 49.06% 47.35% 46.94% 46.84% 48.38% 45.93% 47.35%
             Statistics
Min 18.66% 18.78% 19.90% 20.07% 18.13% 10.11% 16.40% 16.15% 12.82% 9.37% 20.45% 19.89%
25% Quantile 38.21% 37.61% 39.25% 40.53% 34.03% 41.41% 40.43% 35.55% 36.12% 40.03% 34.43% 36.69%
Mean 41.79% 41.15% 42.24% 43.19% 41.72% 43.37% 43.27% 42.16% 41.27% 42.80% 40.30% 41.02%
Median 42.89% 41.66% 41.92% 42.91% 41.12% 44.70% 43.64% 41.82% 43.43% 43.90% 40.37% 41.90%
75% Quantile 47.62% 47.79% 47.78% 48.74% 48.34% 49.21% 48.01% 48.67% 46.87% 48.49% 46.43% 47.49%
Max 57.40% 57.35% 58.78% 59.81% 58.03% 59.05% 59.41% 63.08% 56.64% 59.14% 59.77% 57.65%
TABLE V: Results of FDR (in percentage), MT (in min) and TSR (in percentage) of LTM for the $Closure$ project and a $50\%$ minimization budget. The highest FDR and TSR as well as the shortest MT are highlighted in bold.
 Metric Approach CodeBERT GraphCodeBERT UniXcoder StarEncoder CodeLlama-7b
Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Euclidean Cosine Euclidean
FDR 0.77 0.77 0.78 0.78 0.79 0.74 0.73 0.76 0.77 0.76
MT 15.02 30.53 19.73 25.98 17.80 27.61 16.45 34.17 54.87 66.84
TSR 53.18% 51.98% 53.75% 52.53% 52.55% 54.27% 54.40% 52.51% 53.81% 53.70%
Achieved Saving in Testing Time. Table IV shows that, for the $50\%$ minimization budget, LTM achieved an average TSR ranging from $41.15\%$ to $43.37\%$, across configurations. The average TSR achieved by UniXcoder/Cosine was $41.72\%$. This indicates that, reducing the number of test cases by $50\%$ results in a $41.72\%$ reduction in testing time.
RQ1.2 results. Table III shows that the average MT of all LTM configurations ranges from $0.61$ to $14.74$ min per project version. The MT of LTM using UniXcoder/Cosine, which achieved the highest average FDR ($0.84$), also appears to be efficient with $mean~{}=~{}0.82~{}min$ and $median~{}=~{}0.35~{}min$.
Preparation Time. The average preparation time ranges from $0.26$ to $17.71$ min per project version, across configurations. We observed that the preparation time of UniXcoder/Cosine is $65.43$ times shorter than CodeLlama-7b. This is because (1) larger language models require much longer model loading time ($12.07$ min for CodeLlama-7b, compared to UniXcoder for $0.04$ min, on average) and significantly larger memory ($13.10GB$ for CodeLlama-7b compared to $504MB$ for UniXcoder) and (2) the embeddings generation time of larger language models is significantly longer ($5.15$ min for CodeLlama-7b compared to $0.16$ min for UniXcoder, on average) as they have more complex architectures and a larger number of parameters, making the generation of embeddings more time-consuming.
Search Time. The average search time of LTM ranges from $0.46$ to $1.90$ min per project version, across configurations, with a maximum of $24.49$ min. LTM using CodeBERT with Cosine similarity was the fastest in terms of search time, ranging from $0.04$ to $3.86$ min, across project versions. Not only LTM using UniXcoder/Cosine took a shorter search time ($0.78$ min) than the average search time across configurations ($1.03$ min), it also achieved the highest FDR among all configurations, thus making it the best option for LTM.
Results for the Closure project. We further evaluated LTM on the $Closure$ project, which has $174$ versions with an average of $7,308$ test cases per version. Though this project is part of the Defects4J dataset, it was previously excluded from the ATM evaluation due its scalability limitations [8] and the large size of our experiments. However, evaluating the effectiveness and efficiency of LTM on this project, considering its larger scale and test suite, can be informative. Table V presents the results regarding FDR, MT (in min), and TSR (in percentage), respectively. We can see that UniXcoder/Cosine achieves once again the highest FDR of $0.79$ and an MT of only $17.90~{}min$, while saving $52.55\%$ testing time, which indicates that LTM can be effective and scalable for larger projects and test suites.
RQ1 summary. LTM achieved high FDR results ($0.79$, on average, across configurations) for a $50\%$ minimization budget. UniXcoder/Cosine is the best LTM configuration when considering both effectiveness ($0.84$ FDR on average) and efficiency ($0.82$ min on average), with an average TSR of $41.72\%$
4.3.4 RQ2 results
RQ2.1 results. For a $50\%$ minimization budget, Tables II, III, and IV report FDR, MT, and TSR, respectively, for the best configuration of LTM (i.e., UniXcoder/Cosine) and the two best configurations of the baseline approach: ATM with combined similarity and tree edit distance similarity.
We observe that LTM significantly outperforms ATM/Combined, the best ATM alternative, and ATM/TreeEditDistance in terms of FDR, with $\alpha=0.01$. Compared to ATM, LTM achieves higher average FDR ($0.84$) with lower standard deviation ($0.06$ compared to $0.10$ for ATM/Combined and $0.08$ for ATM/TreeEditDistance).
Figure 3 shows that UniXcoder/Cosine converges faster to a higher FDR ($0.84$ across projects), compared to ATM/Combined ($0.80$ across projects) and ATM/TreeEditDistance ($0.81$ across projects). This may be explained by the fact that (1) UniXcoder utilizes both code comments and AST information from the source code, whereas ATM ignores code comments, and (2) UniXcoder was pre-trained on various code understanding tasks to help the model learn the contextual information for each token, and thus the meaning of each code element (e.g., method, variable names, and string characters) in the input test code. In contrast, ATM uses tree-based similarity based on ASTs, which lack information regarding the syntactic context and certain details of the test code.
For some projects, ATM achieved higher FDR than LTM. However, a Fisher’s Exact Test suggests that, for these projects, there is no significant difference between the FDR results of ATM and LTM. Nevertheless, a possible explanation for such performance difference could be that the percentage of test code exceeding the token limits of language models on these projects (5.72%, 3.21%, 3.25%, and 7.25% for $Chart$, $Compress$, $Collections$, and $Math$, respectively) is higher than the average of 2.98% across projects, and that the truncated parts are important for distinguishing test cases. Another possible explanation could be due to variability across test cases in some projects (e.g, $Cli$ and $Time$), regarding their low-level code elements such as complex string manipulation and domain-specific character strings, rather than structural information. This makes tree-based similarity better at capturing these differences, by comparing AST node labels, than using the Cosine Similarity of embeddings.
Note that the way ATM preprocessed test cases, by removing all the comments and assertions, may also influence its performance. For example, for the $Codec$ project, 26.18% of the test methods contain comments, which is higher than the average percentage across projects (14.85%). Such comments may convey rich information about test code, such as test case descriptions and expected results. Therefore, due to such loss of information in code comments, ATM may not efficiently achieve higher FDR for this project.
The average number of generations required by LTM was $63.10$ across project versions (maximum = $100$), which is lower than ATM/Combined (mean = $85.60$ and maximum = $130$) and ATM/TreeEditDistance (mean = $74.88$ and maximum = $115$). This indicates that LTM converges faster across generations than ATM, in terms of fitness value, which is the sum of the maximum squared similarity values for all test case pairs per version. Related to this, if due to time constraints a fixed number of generations is used as the termination criterion for search, say $40$ generations, it is clear based on Figure 3 that UniXcoder/Cosine would yield the highest FDR. This indicates that LTM offers better guidance for the GA search, making it more effective at removing redundant test cases while maintaining a higher fault detection rate, thus explaining the results in Table II and why LTM is a better option than ATM for test suite minimization in practice.
Achieved Saving in Testing Time. The average time saving rate of the best configuration of LTM ranges from $18.13\%$ to $58.03\%$ across projects, with a mean of $41.72\%$ across projects. We observe that this saving rate is slightly higher than ATM/Combined ($40.03\%$) and ATM/TreeEditDistance ($41.02\%$).
RQ2.2 results.
Figure 4 depicts MT, preparation time, and search time—the first one including the last two—as a function of the number of test cases per version, for the $661$ versions of the $16$ projects. This is done for the best configuration of LTM (i.e., UniXcoder/Cosine) and compared to the best configurations of ATM (ATM/Combined and ATM/TreeEditDistance) for the $50\%$ minimization budget.
Figure 4: Scatter plots of the number of test cases and MT, preparation time, and search time (in min), for LTM (UniXcoder/Cosine) and ATM, across all the $661$ project versions for the 50% minimization budget
We observe that LTM ran much faster than ATM in terms of MT ($79.80\%$ and $97.72\%$ faster than ATM/Combined and ATM/TreeEditDistance on average, respectively) with both much less preparation time ($92.70\%$ and $99.35\%$ less than ATM/Combined and ATM/TreeEditDistance on average, respectively) and search time ($35.16\%$ and $7.81\%$ less than ATM/Combined and ATM/TreeEditDistance on average, respectively). In addition, we observe that the observations in Figure 4 follow a quadratic relationship rather than a linear one. Therefore, a quadratic regression model is fitted for the preparation time, search time, and MT as a function of the number of test cases per version, respectively, to quantify and compare relationships across approaches.
The equation of the quadratic regression model is as follows.
$Time=a*n^{2}+b*n+c$ (8)
where $n$ denotes the number of test cases per project version, $a$ and $b$ are regression coefficients, and $c$ denotes the intercept of the quadratic regression model.
The results of quadratic regression confirmed that the MT and the number of test cases, for both LTM and ATM, follow a quadratic relationship as the coefficients on the quadratic term were statistically significant with $\alpha=0.01$. This is, as expected, due to the fact that (1) similarity values are measured between test case pairs, which increases quadratically with the number of test cases, and (2) the search space for GA increases rapidly with the number of test cases. The $R^{2}$ values for the regression model of LTM and ATM/Combined were $98.89\%$ and $89.51\%$, respectively. This indicates that a high percentage of the variation in MT can be explained by the number of test cases. However, the $R^{2}$ value for ATM/TreeEditDistance was only $50.36\%$, indicating that a much smaller proportion of the variation in MT is be explained by the number of test cases in this case. This is because the preparation time of ATM/TreeEditDistance, which is part of its MT, was also largely influenced by AST sizes. More discussions will be provided in the following section.
Moreover, a comparison of regression coefficients indicates that the MT of LTM increases up to two orders of magnitude slower than ATM with the squared number of test cases per project version: $8.255e^{-06}$ for LTM versus $9.314e^{-05}$ and $4.446e^{-04}$ for ATM/Combined and ATM/TreeEditDistance, respectively. For large industry projects with large test suites, LTM is thus much more scalable in terms of MT, and therefore a better option than ATM in practice.
Preparation Time. We found that LTM is much more efficient, in terms of preparation time across project versions (mean = $0.23$ min and median = $0.14$ min), compared to ATM/Combined (mean = $3.15$ min and median = $0.74$ min) and ATM/TreeEditDistance (mean = $35.31$ min and median = $6.10$ min). Moreover, the preparation time of LTM takes only $35.75\%$ of the MT per version, that is much less than ATM/Combined ($71.22\%$) and ATM/TreeEditDistance ($93.74\%$). Quadratic regression shows that the coefficient on the quadratic term for LTM ($1.241e^{-06}$) is one order of magnitude smaller than the coefficients for ATM/Combined ($6.634e^{-05}$) and two orders of magnitude smaller than ATM/TreeEditDistance ($4.3680e^{-04}$), with all coefficients statistically significant with $\alpha=0.01$. This indicates that, on larger industry projects with much larger test suites, the absolute calculation time difference between LTM and ATM is expected to be extremely large, thus making LTM a better option regarding scalability in preparation time.
The $R^{2}$ values for the regression model of LTM and ATM/Combined are $97.70\%$ and $87.86\%$, respectively. This indicates that more than $87\%$ of the variation in preparation time can be explained by the number of test cases. However, the $R^{2}$ value for ATM/TreeEditDistance is $49.96\%$, which is mainly due to its preparation time as it is highly influenced by test code length. For instance, as shown in Figure 4, there are project versions with nearly $4k$ test cases from the $Time$ project, which have less preparation time than project versions with around $3k$ test case from the $Math$ project. This is due to that fact that the average test code for the $Math$ project is longer than that for the $Time$ project, resulting in its AST tree size being twice as long. Therefore, although the test suite size of the $Math$ project is smaller, its preparation time is significantly longer than that of the $Time$ project. In contrast, the preparation time for LTM is not affected by the length of test code, which further supports our claims regarding its scalability.
Search Time. We observe that the time LTM takes for searching the optimal test suite (mean = $0.59$ min and median = $0.21$ min across projects) is also shorter than that for ATM/Combined (mean = $0.91$ min and median = $0.29$ min) and ATM/TreeEditDistance (mean = $0.64$ min and median = $0.23$).
Moreover, quadratic regression reveals that, while quadratically increasing with the number of test cases per version, the search time of LTM increases more slowly with the squared number of test cases per version (a coefficient on the quadratic term equal to $7.013e^{-06}$) than ATM/TreeEditDistance ($7.844e^{-06}$) and one order of magnitude more slowly than ATM/Combined ($2.680e^{-05}$), with more than $79\%$ ($R^{2}$) of the variation in MT being explained by the number of test cases.
Note that the search time not only depends on the number of test cases per version but is also impacted by the number of generations required by the GA to converge. For instance, though the number of test cases per version for the $Math$ project ($2,078$) is less than that for the $Time$ project ($3,918$), GA needed more generations to converge for $Math$, thus resulting in longer search time. Test cases of the $Math$ project are more diverse than that of the $Time$ project, with a higher standard deviation of similarity values ($0.13$ compared to $0.09$), thus leading to more generations and longer convergence time for the GA search.
In conclusion, in addition to being more efficient than ATM in terms of preparation time, the similarity measure employed by LTM provides better guidance and thus helps the search algorithm converge faster than ATM, thus resulting in less search time.
RQ2 summary. LTM outperforms ATM by achieving significantly higher FDR ($0.84$ versus $0.81$, on average). However, the main benefit of LTM is that it is running much faster than ATM ($0.82$ min versus $4.06$ min, on average), in terms of both preparation time (up to two orders of magnitude faster) and search time (up to one order of magnitude faster). Given the acute scalability issues met in industrial contexts regarding minimization for large test suites and systems, this result is of high practical importance.
4.4 Discussion
4.4.1 Scalable test suite minimization with much less preparation time than the SOTA.
Our results show that LTM runs up to two orders of magnitude faster than the SOTA approach (i.e., ATM) while achieving a statistically significantly higher fault detection rate. We should note that, according to the average MT across projects, LTM runs five times faster than ATM/Combined and $44$ times faster than ATM/TreeEditDistance. These differences tend to increase significantly as the test suite size grows. For example, for the $Math$ project with $2078$ test cases, LTM runs $6$ and $101$ times faster than ATM/Combined and ATM/TreeEditDistance, respectively. This suggests that LTM, while achieving higher effectiveness, can save enormous time and resources for minimizing test suites of the much larger software systems and test suites encountered in practice. The preparation phase plays a crucial role in the scalability of LTM. Since we employed pre-trained language models to extract vector-based test method embeddings and used a highly optimized function to compute similarity between them, we managed to achieve much shorter preparation times compared to ATM. In addition, unlike ATM, the preparation time of LTM is not impacted by test code length, which further supports our claims about its scalability. Furthermore, since we used the test code as input without preprocessing it and converting it into other formats, such as ASTs for ATM, we saved substantial memory space as well.
Unlike coverage-based techniques, the size of a project (i.e., source lines of code) has no impact on the minimization time of black-box TSM approaches, including LTM and ATM, as minimization time mostly depends on the number of test cases per project version. In addition, we optimize the search process by changing the data structure for input similarity values and accelerating the fitness calculation, which reduced the search time by $272.90$ times. The $Closure$ project—the largest by far, comprising $179$ KLoC and $174$ versions, with $7,308$ test cases per version—took only $17.80$ min in terms of MT and achieved a FDR of $0.79$, for a 50% minimization budget. Note that for ATM, $Closure$ was far too large to even run our experiments, thus further illustrating the scalability gains with LTM.
Though LTM uses the same search termination criterion as ATM, which is defined as a fitness improvement of less than $0.0025$ across generations, we observed that FDR reaches a plateau after a certain number of generations (e.g., about an average of $65$ generations across project versions using LTM with UniXcoder/Cosine). However, the GA search continues as long as there is sufficient improvement in fitness (e.g., LTM with UniXcoder/Cosine reached a maximum of $100$ generations). This suggests that the search termination criterion unnecessarily prolongs the search and thus needs further investigation. For example, using a fixed number of generations, determined based on experience with past versions, could save time and resources while achieving high FDR values.
4.4.2 Effective test suite minimization reduces testing cost.
Further, LTM achieves an average saving of $41.72\%$ in test execution time with a $50\%$ test suite minimization budget. However, though this is due to the relatively modest size of the systems we experiment with, one might point out that the average execution time of test suites in our dataset is smaller than LTM minimization time. Nevertheless, LTM still offers significant advantages as testing is performed frequently, especially in CI contexts, while minimization is occasional, unlike test case selection and prioritization. It can thus achieve substantial savings in time and resources over time, particularly in CI environments, where numerous builds are tested for every code change, thus warranting the elimination of redundant or unnecessary test cases to significantly reduce the overall test execution time and required testing resources (e.g., memory, CPU).
4.4.3 More informative similarity measures that better guide the search.
Our results show that the similarity measures LTM employed better guide the GA to converge faster to a higher FDR, when compared to ATM, confirming the usefulness of using pre-trained language models for extracting test method embeddings.
Results also suggest that the FDR results are influenced by multiple factors including the language models, distance functions, and test code characteristics. Language models differ in the types of input used for pre-training, the model architectures employed, and the kinds of tasks they were pre-trained for, thus generating embeddings with various characteristics. While Cosine Similarity captures directional similarity by measuring the angle between embeddings [48], Euclidean Distance captures magnitude similarity by measuring the straight-line distance between embeddings in a multidimensional space [43]. Therefore, using different combinations of language models and distance functions is expected to yield varying FDR results.
In addition, the test code characteristic (e.g., domain-specific character strings or complex string manipulations) of certain projects may also impact FDR performance, depending on the input (test code or ASTs) and dataset (size and diversity of programming languages ) used for pre-training the language models. In conclusion, the performance of LTM in terms of FDR may be affected by several confounding factors. Future research should explore the effects of these factors through deeper analyses of test code structures and controlled experiments.
4.4.4 Language models with fewer parameters may offer a better trade-off than those with many more parameters in TSM contexts.
The results show that UniXcoder ($125Million$ parameter size) outperforms CodeLlama-7b ($7Billion$ parameter size) in terms of both effectiveness and efficiency. This suggests that larger language models, though comprising substantially more parameters and being pre-trained on larger datasets, may not necessarily perform better for certain downstream tasks such as TSM. CodeLlama-7b, as a decoder-only model, was pre-trained for code generation tasks, such as code infilling. While generating (infilling) the next token, the structure of the decoder makes each token only learn information from preceding tokens, thereby potentially making its embedding less informative due to the absence of information from subsequent tokens. In contrast, code understanding tasks (e.g., MLM) allow the model to learn the information before and after each token, thus making it more effective in learning the contextual semantic information from the source code. Furthermore, UniXcoder enhanced the code representation by utilizing AST as input, which further enables the model to understand the source code structure and grammar, thus resulting in better FDR.
Note that, to the best of our knowledge, there is no publicly available open-source language model that outperforms UniXcoder on the code clone detection task, which is similar in some ways to our minimization task. More importantly, given that scalability is crucial for LTM, larger language models may require much more memory, computational resources, and time when generating embeddings due to their substantially higher number of parameters and complex model architectures. Therefore, for a downstream task, such as TSM, that requires assessing test code similarity, relatively smaller language models can be a better option than larger ones.
4.4.5 Using full code versus preprocessed code.
LTM used the source code of test cases without preprocessing as input to the language models and yielded better results than using preprocessed code, as shown in our preliminary study. However, some preprocessing techniques are still worth investigating, such as normalizing variable identifiers, since they might not have an impact on the data flow in the source code. Therefore, using such techniques may potentially help language models generate embeddings that are more generalizable and suitable for code similarity calculation across projects, thus improving the fault detection power of LTM.
5 Threats to validity
Internal Validity Internal threats to validity are concerned with the ability to draw conclusions from our experimental results. To make fair comparisons and draw valid conclusions, we conducted our experiments using the same experimental design and dataset used to evaluate ATM [8]. We evaluated the alternative LTM configurations and compared the best configuration to the two best configurations of ATM. For minimizing test suites, we used the same fitness function and parameter settings as ATM, but we further optimized GA by utilizing a more efficient data structure for storing the input similarity values and accelerating the fitness calculation. We compared the performance using optimized GA for both LTM and ATM. However, ATM used preprocessed code as input, whereas LTM used source code without preprocessing. We did not consider using the full code for ATM due to its scalability issues. Another threat is related to the token length limit of language models, which might truncate the source code of some test cases [49], thus resulting in information loss that might in turn negatively impact the performance of LTM. However, we observed that, for each version, the percentage of test methods with the token length exceeding the limit is only $2.98\%$ across projects, which can be considered negligible. Nevertheless, some projects have a relatively higher percentage of test methods exceeding that limit (e.g., $7.25\%$ for $Math$), which might explain the slightly lower FDR of LTM compared to ATM on these projects. Future research should further investigate the use of additional language models with longer token length limits, though they might greatly impact efficiency.
External Validity External threats are concerned with the ability to generalize our results. We assessed both the performance of ATM and LTM on a large dataset consisting of $16$ Java projects with a total of $661$ versions collected from Defects4J for which we could trace failures to faults. We further validated the performance of LTM using an additional larger project from Defects4J, called $Closure$, consisting of $174$ versions, for which we could not evaluate ATM. Though we did not evaluate LTM on projects written in other programming languages, LTM can be easily adapted to other programming languages as the language models we used were pre-trained on multiple programming languages and have been demonstrated to yield good performance in various downstream tasks in other programming languages [10, 11, 12, 34]. Future research should extend our study to investigate LTM on projects with other programming languages to further generalize our conclusions. In addition, test method execution time may vary from run to run. To measure the achieved time-saving in test execution time more precisely, we used the Defects4J infrastructure to run test suites $10$ times and collect the average execution time for each test method in a clean, stable environment, which ensures we run each test suite without interference and fluctuations in memory or CPU usage (e.g., exclusive access to a node).  [^8] We further collected the minimum and maximum execution times for each test method across the $10$ runs. The results show that the difference in TSR, based on minimum and maximum execution times, ranges from 0.05% to 1.38% across all configurations, which indicates that the variation in test method execution times across runs is small and does not impact our conclusions regarding TSR. However, we found many test cases ($24.14\%$ across projects) with 00 execution time as they took less than $0.0001$ seconds to run.  [^9] To ensure this did not impact our results, we investigated the effect of these test cases on the Time Saving Rate by assuming their execution times to be $0.0001$—their maximum possible value—and re-calculating the Time Saving Rate. We found that the average Time Saving Rate of all configurations slightly increased by 0.08% to 0.09% only, which is negligible and thus does not invalidate our conclusions. Future research should consider evaluating TSM approaches on much larger industrial systems.
6 Conclusion
In this paper, we proposed LTM, a scalable and black-box similarity-based test suite minimization (TSM) approach based on pre-trained language models. We investigated five alternative language models—CodeBERT, GraphCodeBERT, UniXcoder, StarEncoder, and CodeLlama—to extract test method embeddings and two similarity measures (i.e., Cosine Similarity and Euclidean Distance) for similarity calculation based on these embeddings. We employed an optimized version of Genetic Algorithm (GA) by utilizing an efficient data structure for handling similarity values to improve fitness calculation and finding optimal subsets of test suites. We assessed the performance of alternative LTM configurations on a large dataset consisting of $17$ Java projects with a total of $835$ versions collected from Defects4J. We used the Fault Detection Rate (FDR), Minimization Time (MT), and Time Saving Rate (TSR) as evaluation metrics to assess the effectiveness and efficiency of LTM. We identified the best configuration of LTM and compared it to ATM, a recent black-box TSM approach that achieved better trade-off in terms of effectiveness and efficiency than the alternative SOTA TSM approach (FAST-R). Results indicate that the best LTM configuration (i.e., UniXcoder/Cosine) outperformed the two best configurations of ATM by achieving significantly higher FDR results ($0.84$ versus $0.81$, on average) and more importantly, running much faster ($0.82$ min versus $4.06$ min, on average) than ATM, in terms of both preparation time (up to two orders of magnitude faster) and search time (one order of magnitude faster), which is particularly important on much larger industrial systems and test suites, while achieving a saving of $41.72\%$ in test execution time for a $50\%$ test minimization budget.
Future work. We plan to assess the performance of LTM in practice using large industrial software systems and projects in other programming languages to further generalize our conclusions. The main challenge is the ability to automatically trace test failures to system faults.
7 Data Availability
Data, detailed results, and implementations will be made available upon the paper’s acceptance.